OpenStudy (anonymous):

Write the equation of a line in point-slope form that is parallel to the line y = 3x +6 and passes through the point (-5, 8)

OpenStudy (anonymous):

y - 8 = 3(x + 5) y = 3x + 15 + 8 y = 3x + 23